If the intensity of sound is doubled, by how many decibels does the sound level increase?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. We know that, β = 10 log base 10 J/I base 0 According to the questions β base A = 10 log base 10 (2I/I base 0) ⇒β base B – β base A = 10 log (2I/I) = 10 × 0.3010 = 3 dB.